At an activity level of 10,000 units, total variable costs were $35,000 while total fixed costs were $20,800. If 16,000 units are produced and this activity is within the relevant range, which of the following statements is correct?
Answer #1

(Q)  At an activity level of 10,000 units, total variable costs were $35,000 while total fixed costs were $20,800. If 16,000 units are produced and this activity is within the relevant range, which of the following statements is correct?

(Ans) total unit cost would equal $4.80

  • At 16,000 units, FC stays $20,800, so FC/unit = $20,800 / 16,000 units = $1.30, VC/unit = $35,000 / 10,000 units = $3.50. At 16,000 units, total VC = 16,000 x $3.50 = $56,000. Total cost = VC + FC = $56,000 + $20,800 = $76,800
  • Cost per unit = $76,800 / 16000 = $4.80.